Sol Summary – March 19th

Summary Title : Day #2 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary: XO arrived at Hab with the remainder of our research equipment. Crew 207 set up 3D printing systems in Level 1 of the Hab and RAM and set up the bike-powered 3D printing apparatus in Level 1 of the Hab. COM prepared the second part of thoracentesis project in the Science Dome. ENG checked the status of EVA suits. XO provided a fresh battery for the RAM smoke detector. Weather was clear and sunny for EVA #2 to Cattle Gate and Zubrin’s Head. XO and GH planned tomorrow’s EVA. AST continued operations in the solar observatory.

SCI and HSO 3D printed the following labware items out of
biodegradable, food-safe plastic:

Microscope cover slide

Western blot comb

Soil sieve

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Shade cover of mobile devices (including iPad, iPhone + GPS)

Look ahead plan : We will conduct a walking EVA to scout out a location suitable for drone mapping of MDRS. We will continue to 3D print items at MDRS to save time and money for future crews. We will initiate data collection for our thoracentesis and bike-powered 3D printing projects.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA : Cattle Gate, Zubrin’s Head

Reports to be filed : Operations, GreenHab, Journalist, EVA Report, EVA Request, Astronomer

Support Requested : None

EVA Report – March 19th

Crew 207 EVA Report 19-Mar-2019

EVA #2

Author: Tiffany Ni

Purpose of EVA: The goal of this EVA was to reach the highest point near Cow Dung Road that directly leads into MDRS from the highway turnoff to help us conduct drone mapping and 3D printing a
high-resolution map.

Start time: 1410h

End time: 1600h

Narrative: After a 5 minute depressurization, our EVA team (Julielynn, Kevin, and Tiffany) checked the state of the rovers – they were charged and ready. Kevin and Tiffany drove Curiosity while Julielynn drove Spirit. Our departure from the Hab was at 1410h. We unplugged our rovers and headed out along Entrance Road and turned onto Cow Dung Road. We continued along Cow Dung Road, passing Kissing Camel Ridge and Zubrin’s Head. We parked the rovers along the right side of Cow Dung Road – 100m away from the Cattle Guard – and began our hike. The hills around Cattle Guard did not appear safe for a hike on foot so we decided to walk to the fence instead. Here, we used the GPS device to determine the exact coordinates for our flight planning software for drone mapping (0520349, 4247084). This is far more precise than the approximated coordinates we initially determined from the map of MDRS. Next, we turned the rovers around and drove along Cow Dung Road. We stopped at Zubrin’s Head and again, using our GPS device, we determined the exact coordinates (0519362, 4248196). We arrived back at the Hab at 1600h with 97% battery for Curiosity (84.4h) and 66% battery for Spirit (85.1h).

Destination: Cattle Guard

Coordinates: Cattle Guard (0520349, 4247084), Zubrin’s Head (0519362, 4248196)

Participants: Julielynn Wong (Comms, EVA Leader), Kevin Ho (HSO), Tiffany Ni (SCI)

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Driving and Walking

Operations Report – March 19th

Crew 207 Operations Report 19 Mar 2019

SOL: 3

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 11:00 hours

Generator was turned off at 07:30, SOC 100%

Diesel Reading – 54 %

Propane Reading – 68 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 290 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 170 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 141904.7 gallons

Toilet tank emptied: Yes

Deimos rover used: No, not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not applicable
Beginning charge: Not applicable
Ending charge: Not applicable
Currently charging: Not applicable

Spirit rover used: Used
Hours: 85.1 hours
Beginning charge: 100 %
Ending charge: 66 %
Currently charging: Yes

Opportunity rover used: Not used
Hours:53.6 hours
Beginning charge: 100 %
Ending charge: 56 %
Currently charging: Yes

Curiosity rover used: Used
Hours: 84.4 hours
Beginning charge: 100 %
Ending charge: 97 %
Currently charging: Yes

Notes on rovers: EVA mission 2.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where?

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Musk Observatory: Please see Astronomer Report regarding video capture issues.

Questions, concerns and requests to Mission Support: None

Greenhab Report – March 19th

Crew 207 GreenHab Report – 19 Mar 2019

GreenHab Officer: Diane Rothberg

Environmental control: Heat and cooling w/ambient air (9 hr)

Shade Cloth (40%)

Average temperatures:

Low: 16.0 deg C
High: 28.5 deg C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gallons

Water in Blue Tank: 170 gallons

Crops watered at 08:40 hours (3gal) and 17:45 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 08:40 AM 25 deg C / 20%
2) 12:30 AM 25 deg C / 20%
3) 15:05 PM 28 deg C / 20%
4) 17:45 PM 24 deg C / 20%

Changes to crops: None

Narrative:
Harvest: sage 2 gm, parsley 6 gm

Support/supplies needed: None

Journalist Report – March 19th

Journalists Report March 19

At 3:00 am, Dean arrived with all kinds of electronic and mechanical goodies.

After allowing him a short nap, 3 of us assembled a bicycle and mounted it to a friction trainer. Dean hooked up a battery and ancillary monitoring devices. Each crew members will ride for an hour per day for at least 2 days. Energy will be stored in the battery, and used to power the 3D printers. This activity keeps us in shape, maintains high crew morale, and creates electrical power
independently.

Kevin hooked up a 3D printer in the RAM and printed a microscope slide holder, a soil sieve base, and a mold for making protein gels (used for Western blotting). How wonderful that these aids to science can be made right on Mars without the necessity of transporting them up here!

In the afternoon, Kevin, Tiffany and Julielynn went on an EVA to scout the turn off the main highway to Cow Dung Road. As there is no signage, they plan to drone map the area to create a 3D printed map which might be a helpful aid to pinpointing the intersection.

In the evening, Julielynn trained all crew members in needle decompression skills to treat a potentially fatal tension
pneumothorax. She employed her leading-edge 3D printed medical simulation training devices. While this was fascinating and exciting it was kind of scary for those of us who are not in the medical profession! But being on Mars, where the nearest hospital is months away, it’s a handy skill to know!

Cheers,

Julielynn Wong, MD, MPH, FACPM
Chief Medical Maker

Astronomy Report – March 19th

Astronomy Report

Name: Tom Baldwin

Crew: 207

Date: 19-MAR-2019

MDRS ROBOTIC OBSERVATORY

Robotic Telescope MDRS-WF: downloaded captured images of Horsehead Nebula

Objects Viewed: Horsehead Nebula

Problems Encountered: none

Processed Crab Nebula photo, attached JPEG

MUSK OBSERVATORY : Viewed Sun through eyepiece and camera

Solar Features Observed: Sun

Problems Encountered: Captured images did not download into Solar Image folder. Steps listed on pages 21 through 23 followed.